OpenStudy (anonymous):

If a sphere has a radius of 2 cm, what is the approximate surface area of the sphere?

OpenStudy (anonymous):

formula :: \[4 \pi r ^{2}\]

OpenStudy (anonymous):

now replace r by the value of the radius and you will get the answer

OpenStudy (anonymous):

idk how to

OpenStudy (anonymous):

answer ? @yajna

OpenStudy (anonymous):

that's not a choice

OpenStudy (anonymous):

@superman36

OpenStudy (anonymous):

what are the options ?

OpenStudy (anonymous):

12.1 33.5 50.3 100.5

OpenStudy (anonymous):

answer is 50.3

OpenStudy (anonymous):

\[4 \times \pi \times (2)^{2}\] = 50.3